OpenStudy (rainbow_dash):

Y varies directly as x and inversely as z. y=2/9 when x=2 and z=6. Find y when x=3 and z=5. I can solve this on my own if i know the setup. is directly and inversely meaning that the setup is y=x/z?

OpenStudy (phi):

y varies directly as x means y= kx for some number k y varies inversely as z means y= m/z for some number m

OpenStudy (rainbow_dash):

Oh, is K the constant? finding that will allow me to find the new equation, right

OpenStudy (phi):

putting them together it looks like y= kx/z use the given values to find k. now you can find y for other values of x,z

OpenStudy (rainbow_dash):

Okay so it'll be y=kx/z, which is 2/9=K2/6. then 2/9*6=K2, divide by 2 equals K. then I'd take K and plug in to y=K3/5..

OpenStudy (phi):

yes. though I would be tempted to reduce 2/6 to 1/3 because it looks simpler. But either way...

OpenStudy (rainbow_dash):

okay the answer is y = 2/5 in the final result i'm thinking. thanks! it's just the equations/formulas with variation (or any math for that matter) always escape me during the summer.

OpenStudy (phi):

looks good.